OpenStudy (anonymous):

If a dilation has center at the origin and scale factor 2, then what is the preimage point that would map to (4,8)?

OpenStudy (anonymous):

How do you solve this?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you have some unknown point (x,y)

jimthompson5910 (jim_thompson5910):

x maps to 4 y maps to 8

jimthompson5910 (jim_thompson5910):

k*x = 4 2x = 4 .. the scale factor k is 2 x = ??

OpenStudy (anonymous):

X would be 2

OpenStudy (anonymous):

So you would just divdie by two.

jimthompson5910 (jim_thompson5910):

k*y = 8 2*y = 8 y = ??

OpenStudy (anonymous):

4

OpenStudy (anonymous):

So the answer would be (2,4)

OpenStudy (anonymous):

So lets day if you have another question saying that the scale factor is three and the preimage points are (3,12) Would you divide by 3?

jimthompson5910 (jim_thompson5910):

look at this drawing http://www.winpossible.com/App_Themes/default/Images/CourseImages/Transformation.JPG

jimthompson5910 (jim_thompson5910):

Notice how the preimage is the original and the image is the new figure (after you do some transformation to it)

jimthompson5910 (jim_thompson5910):

in the previous problem, we had the image (4,8) we didn't know the preimage, so we undid multiplication to find it

OpenStudy (anonymous):

Oh okay I get it so the first answer would be (2,4)

jimthompson5910 (jim_thompson5910):